OpenStudy (anonymous):

One number is 3 less than twice another. If their product is 405, find the numbers. Can I get some help on this?

OpenStudy (anonymous):

been working with quadratic equations

OpenStudy (anonymous):



OpenStudy (anonymous):

I am not sure , but thought that if a word problem says the product, doesn't that mean to multiply?

OpenStudy (anonymous):

oopsy i feel dumb. so change the second one to xy=405

OpenStudy (anonymous):

2x + x - 3 = 405 3x+ 408 x= 272 x-3 = 133

OpenStudy (anonymous):

272+133=405

OpenStudy (anonymous):

Yeeha!!

OpenStudy (anonymous):

Thanks, you helped me think of it in a different approach. That jarred my memory